OpenStudy (anonymous):

what is the y-intercept of the line that passes through the points (3,7) and (-2,-3)?

OpenStudy (anonymous):

it crosses the y intercept at 1

OpenStudy (anonymous):

first find the slope slope = (y2-y1)(x2-x1)

OpenStudy (anonymous):

then use point slope form to get the equation of line from that equation you will get y-intercept point slope form is y-y1 = m(x-x1)

OpenStudy (hba):

After you get the equation of the line At y-intercept x=0 You get y-intercept :)